solve the solution
y=3^(2x+1)
y=9^(4-2x)
do you make the powers = one another for 2x+1=4-2x

Found 2 solutions by vleith, stanbon:
Answer by vleith(2983)

Almost. just remember that 9 = 3^2
y=9^(4-2x)
y =(3^2)^(4-2x)
y = 3^(8-4x)
now set them equal and solve

Answer by stanbon(75887)

To equate the powers you have to have equal bases:
y=3^(2x+1)
y=9^(4-2x)
------------
y = 3^(2x+1)
y = 3^[2(4-2x)]
-----------------
2x+1 = 2(4-2x)
2x+1 = 8-4x
6x = 7
x = 7/6
===========
Cheers,
Stan H.